# Price Experimentation and Interference

### v4

@@ 7.4 Properties of Biases with Market Balance
Below we provide two figures to complement our limit results. In Figures 1 and 2 we present bias normalized by the arrival rate $\lambda$. We normalize by $\lambda$ instead of the GTE because the GTE for profits converges to 0 around the optimal price. As $\lambda$ increases, the market becomes more supply-constrained and bias for a LR experiment decreases as competition between listings dissipates. Figure 2 shows the normalized bias of demand and profits in a CR experiment and this dependence on $\lambda$ is naturally reversed. Both figures also show interesting behavior of biases with respect to the baseline price $p$, a novel aspect of our setting relative to Johari et al., (2022). As the price $p$ changes, the sensitivity of demand to price changes as well. In particular, in both CR and LR experiments, both $D_{0y}$ and $D_{1x}$ tend to be smaller at larger prices (i.e., demand becomes more inelastic as the price becomes large); this behavior drives the shape of the bias curve with respect to price, and subsequently profit. Note in particular that LR experiments exhibit interesting behavior when prices are large: despite the fact that we nominally expect bias to increase when the market is demand constrained, and higher prices reduce demand in the market, nevertheless the bias drops when prices are high. This is because of the fact that elasticities change with price, and illustrates the nuanced nature of interference in price experiments.
@@ 7.5 Change of Sign
[…]
@@ 7.5.1 Numerical Results
Figure 3: Illustration of the change of sign region (purple) in an LR-experiment (left) and CR-experiment (right), respectively, with $v(p)=e^{V-p}$, $V=5$, $c=0.5$, $\rho=1$, and $\epsilon=1$. The vertical axis is the (log of) the arrival rate, or market balance, paramter $\lambda$; the horizontal axis is the (baseline) control price $p$, at which the experiment is conducted. The region where condition (a) is not satisfied ($GTE_{\pi}<0$), so there is no change of sign, is in red. The region where condition (b) is not satisfied ($Bias_{\pi}^{LR}<GTE_{\pi}$ or $Bias_{\pi}^{CR}<GTE_{\pi}$, respectively), so there is no change of sign, is in blue.
@@ 7.5.2 Gap from Optimality
One important implication of Figure 3 is the impact that the change of sign region has on the net profits of firms who use canonical price experimentation methods. Consider the practical importance of the change of sign region’s borders. The right border of the change of sign region (where purple meets red), represents the optimal price for that particular market balance parameter $\lambda$. This is the point where $GTE_{\pi}=0$. The left border of the change of sign region (where blue meets purple) represents the final price a firm would reach, were it to repeatedly update it’s price in response to canonical A/B experimentation. This is the point where $\widehat{GTE}_{\pi}=0$, or equivalently, where $Bias_{\pi}=GTE_{\pi}$. Note that this price corresponds to the optimum suggested by repeated canonical A/B experiments. This is because, to the right of that price point, the expriment suggests decreasing the price, and to the left, to increase it.
The difference between these two prices (the true optimal price and the experimentally optimized price) can correspond to significant losses in overall profit. In Figure 4, we see the percent of true optimal profit that a firm will achieve if it uses canonical LR (left) or CR (right) experimentation, for a range of market balance scenarios. What we find is that the profit achieved from optimization via experimentation can be far below the optimal profit, with the impact worsening in highly demand-constrained markets (for LR experiments) or demand-constrained markets (for CR experiments). Even when markets are reasonably balanced (i.e., $\lambda\approx 1$), both LR and CR experimentation will lead to a profit that is more than 4% lower than optimal. Overall, these findings suggest that firms who use canonical A/B methods for price experimentation might be operating well below their potentially optimal profits, even if they are in well balanced markets. Furthermore, these losses can increase significantly as markets become more imbalanced, getting to $20-40\%$ of optimal.
Figure 4: Illustration of the gap from optimal profits that canonical A/B experimentation creates.
@@ 8 An Empirical Case Study
Thus far, we have shown that A/B pricing experiments can get the sign wrong, and that associated profit losses can be significant in a two-sided market model with logit demand. To align our analysis even more closely with practical scenarios, we empirically calibrate our model from the previous section using published data from several Airbnb markets, and investigate the conditions under which the sign-change phenomenon may occur in price experiments within this context. Although our model simplifies the complexities of the Airbnb marketplace in many dimensions, we believe this exercise is valuable because it offers practical insights into how our results can align with real-world experiments. As will be shown in this section, our results suggest that a change of sign is indeed quite plausible in the empirical Airbnb context we study, demonstrating that these errors are not just a knife-edge issue—and thus potentially of high consequence for practitioners.
8.1 Model Calibration
In this section, we calibrate a specific version of our model from Section 7 to several Airbnb markets using the data from Farronato and Fradkin, (2022). Specifically, we consider the model from that section with $v(p)=e^{V-p}$. Note that this is a natural choice implied by a logit random utility model, where utility for the product is quasilinear in the price. For fixed $\varepsilon$, $V$, $\lambda$, and $p$, the resulting market balance equation is:
$\rho-s^{*}(\lambda,p)=\lambda\frac{s^{*}(\lambda,p)e^{V-p}}{\varepsilon+s^{*}(\lambda,p)e^{V-p}}.$ (28)
Recall that then in this model, $D(p)=\rho-s^{*}(\lambda,p)$, and for a given supplier cost $c$, its profit is $\pi(p)=(p-c)D(p)$.
In this exercise, we consider that $GTE$ corresponds to suppliers’ profits and that the experiment is on actual prices. It is worth noting that in practice, Airbnb does not actually set the price of listings, but instead sets fees. However, we have elected to use our price testing framework (in place of the fee-model detailed in Appendix 12) because the estimates made in Farronato and Fradkin, (2022) are better suited to examining price experiments. . Airbnb actively provides pricing guidance to hosts aimed at optimizing their pricing strategies and regularly conducts A/B tests to evaluate these recommendations (Deng and Le, 2023). This setting aligns closely with the price testing framework employed in this section. Furthermore, since fee changes are often passed on to consumers through adjustments to prices, this is a reasonable approach for our analysis.
The goal of our calibration process is to arrive at imputed values of the parameters that define the model: (1) the total mass of listings $\rho$; (2) the outside option value $\varepsilon$; (3) the cost $c$; (4) the value $V$; (5) the arrival rate $\lambda$; and (6) the reference price $p$, from which the platform runs a price experiment.
We now describe how we arrive at these six parameters. We leverage structural estimates made by Farronato and Fradkin, (2022) for Airbnb markets in major U.S. cities. In particular, Farronato and Fradkin, (2022) segment the Airbnb market into four tiers: economy, midscale, upscale, and luxury. They study each of these tiers across 10 cities. We refer to each pair of tier and city as a segment, and use superscript $s$ to denote a segment. For each segment $s$, Farronato and Fradkin, (2022) reports (1) an estimate of the average cost (per night booked), which we denote $\hat{c}^{(s)}$; and (2) an estimate of own price elasticity of demand, which we denote $\hat{e}_{p}^{(s)}$. [^11] These values come directly from tables E8 and E14 in the Appendix of Farronato and Fradkin, (2022).
For our empirical study, we treat each segment as its own instantiation of the market model in (28). For example, if we are currently considering midscale listings in Detroit, then $v(p)$ will represent the average consumer valuation for midscale Detroit listings at price $p$. On its part, $\varepsilon$ will represent the value of the “outside” option relative to this segment, accounting for all other Airbnb segments in that city and the option of not purchasing anything within the Airbnb marketplace (e.g., booking an hotel). Our calibration proceeds according to the following steps:
  1. 1.
    We set the total mass of listings $\rho=1$. This allows us to interpret $s^{*}(\lambda,p)$ as the fraction of listings that are available.
  2. 2.
    As usual, we normalize the value of the outside option, $\varepsilon=1$.
  3. 3.
    We set the cost $c=\hat{c}^{(s)}$.
  4. 4.
    We choose a reference price $p$ at which we want to evaluate price experiments. Actual prices were not reported in Farronato and Fradkin, (2022); in lieu of this information, in our results we vary $p$ in a “reasonable” range, as we describe below.
  5. 5.
    We choose an arrival rate $\lambda$. Again, no proxy for $\lambda$ is reported in Farronato and Fradkin, (2022), and so in our results we also vary $\lambda$ in a “reasonable” range, again as we describe below.
  6. 6.
    Finally, given the chosen values of $\rho,\ \varepsilon,\ p$ and $\lambda$, we choose $V$ so that the demand elasticity $pD^{\prime}(p)/D(p)$ for $D(p)=\rho-s^{*}(\lambda,p)$, with $s^{*}(\lambda,p)$ defined in (28), is equal to $\hat{e}_{p}^{(s)}$ (the estimated elasticity from Farronato and Fradkin, (2022)). It is straightforward to check the price elasticity of demand in our demand model is monotonically increasing in $V$, which guarantees that the fitted value of $V$ is uniquely determined in this approach.
Observe that at the conclusion of this procedure, for each value of $\lambda$ and $p$, we have an imputed demand model. We then use this demand model to determine whether a local CR or LR price experiment yields the wrong sign on the estimate of $GTE_{\pi}$, as discussed throughout the paper.
The two “free parameters” in our calibration are the reference price $p$, and the arrival rate $\lambda$, so we vary them in our subsequent empirical analysis. We vary the reference price $p$ in a plausible range from the cost $c$ to a value above the monopoly optimal price implied by $\hat{c}^{(s)}$ and $\hat{e}_{p}^{(s)}$. In our results we also vary $\lambda$ in a plausible range; in particular, we vary $\lambda$ so that in our fitted model we obtain plausible occupancy rates. The occupancy rate is the fraction of total supply that is booked in equilibrium, defined as $\zeta=1-s^{*}(\lambda,p)/\rho$. We vary $\lambda$ to ensure that we plot occupancy rates from 5% to 95%. As an example, one estimate from external market research on short term rental properties Airdna, (2024) estimates that the average occupancy rate across the entire Airbnb market is approximately 60%, with rates closer to 65% for major cities. This suggests the range of values of $\zeta$ in our results includes realistic values.
In our results, we present our visualizations with the two axes of price $p$ and occupancy rate $\zeta$, the latter in place of the arrival rate $\lambda$. We numerically verified that in this model, each pair of original parameters ($\lambda,p$), yields a unique occupancy pair ($\zeta,p$). Further, we have verified that in all instances of our model, as $\lambda$ increases, the resulting $\zeta$ increases monotonically as well. We make this transformation because the occupancy rate is a more natural metric for interpretation than the arrival rate of customers.
8.2 Results
Figure 5: Illustrations of the change of sign region (purple) in LR-experiments (top row) and CR-experiments (bottom row), respectively, for economy listings (left column) and luxury listings (right column) in Boston. The vertical axis is the imputed occupancy rate, and the horizontal axis is the (baseline) control price around which an experiment is conducted. The region where condition (a) is not satisfied ($GTE_{\pi}<0$), so there is no change of sign, is in red. The region where condition (b) is not satisfied ($Bias_{\pi}^{LR}<GTE_{\pi}$ or $Bias_{\pi}^{CR}<GTE_{\pi}$, respectively), so there is no change of sign, is in blue. We plot prices from the cost estimate $\hat{c}^{(s)}$, to a fixed amount above the critical price $p_{c}^{(s)}$.
Figure 5 shows the results of our model’s estimates for LR (top row) and CR (bottom row) price experiments in Boston. The left two plots show results for the economy tier, while the right two plots show results for the luxury tier. We have chosen to include Boston as an example city, but it is important to note that the trends observed in Figure 5, hold for all segments. (See Appendix 11 for these additional figures for all other segments.)
As in our preceding visualizations, the shaded blue region in each plot is where Condition (a) holds in the fitted demand model, cf. (13), and the shaded red region is where Condition (b) holds, cf. (14). These regions overlap in the purple region where both conditions hold. These are the regions where the estimated treatment effect would have the wrong sign. As noted in the preceding section, our results are plotted with the reference price $p$ on the x-axis (varied in a reasonable range, as described in the previous subsection), and the imputed occupancy rate $\zeta$ on the y-axis.
First, observe that in each figure, the right boundary of the change of sign region is a vertical line at a critical price $p_{c}^{(s)}$. To see why, observe that Condition (a) in (13) holds with equality at price $p_{c}^{(s)}$, with the cost $\hat{c}^{(s)}$ and elasticity $\hat{e}_{p}^{(s)}$ corresponding to the segment $s$. In particular, $p_{c}^{(s)}$ satisfies the first order condition for optimal monopoly pricing given $\hat{c}^{(s)}$ and $\hat{e}_{p}^{(s)}$. Further, this identity is insensitive to the value of the occupancy rate $\zeta$ (or equivalently, the arrival rate $\lambda$). Given the monotonicity in Condition (a), at any higher (resp., lower) values of $p$ than $p_{c}^{(s)}$, Condition (a) is guaranteed to not hold (resp., is guaranteed to hold); and again, these facts are also insensitive to the value of $\zeta$.
More importantly, observe that the figures make clear that changes of sign are prevalent in this model. Notably, across the parameter values we tested, LR experiments in particular seem especially vulnerable to changes of sign. This is consistent with the fact that the observed occupancy rates suggest a relatively demand-constrained market, at least as far as bias in price experiments is concerned. On the other hand, CR experiments exhibit a more nuanced behavior. For such experiments, a change of sign is observed in markets that are highly supply-constrained (e.g., with greater than 80-85% of listings being occupied), and when the reference price is near but still below the critical price $p_{c}^{(s)}$.
It should be noted that although CR price experiments exhibit more favorable behavior in these plots than LR price experiments, LR price experiments are more common in industry. In part this is because platforms are often reticent to offer individual customers different prices for the same product (Phillips, 2021). As a consequence, one potential practical takeaway from our work is that firms may in fact have chosen an experimental design that is more commonly producing the wrong sign in their estimated treatment effects in practice, at least in the settings we have studied. Despite the stylized nature of the model, the prevalence of changes of sign across a wide range of parameters suggests this is a phenomenon of significant consequence for practitioners.

### v5

@@ 7.4 Properties of Biases with Market Balance
Below we provide two figures to complement our limit results. In Figures 1 and 2 we present bias normalized by the arrival rate $\lambda$. We normalize by $\lambda$ instead of the GTE because the GTE for profits converges to 0 around the optimal price. As $\lambda$ increases, the market becomes more supply-constrained and bias for an LR experiment decreases as competition between listings dissipates. Figure 2 shows the normalized bias of demand and profits in a CR experiment and this dependence on $\lambda$ is naturally reversed. Both figures also show interesting behavior of biases with respect to the baseline price $p$, a novel aspect of our setting relative to Johari et al., (2022). As the price $p$ changes, the sensitivity of demand to price changes as well. In particular, in both CR and LR experiments, both $D_{0y}$ and $D_{1x}$ tend to be smaller at larger prices (i.e., demand becomes more inelastic as the price becomes large); this behavior drives the shape of the bias curve with respect to price, and subsequently profit. Note in particular that LR experiments exhibit interesting behavior when prices are large: despite the fact that we nominally expect bias to increase when the market is demand constrained, and higher prices reduce demand in the market, nevertheless the bias drops when prices are high. This is because of the fact that elasticities change with price, and illustrates the nuanced nature of interference in price experiments.
@@ 7.5 Change of Sign
$\text{Condition (a)}\quad GTE_{\pi}\geq 0;\qquad\text{Condition (b)}\quad Bias_{\pi}\geq GTE_{\pi}.$
@@ 7.5.1 Numerical Results
Figure 3: Illustration of the change of sign region (purple) in an LR-experiment (left) and CR-experiment (right), respectively, with $v(p)=e^{V-p}$, $V=5$, $c=0.5$, $\rho=1$, and $\epsilon=1$. The vertical axis is the (log of) the arrival rate, or market balance, parameter $\lambda$; the horizontal axis is the (baseline) control price $p$, at which the experiment is conducted. The region where condition (a) is not satisfied ($GTE_{\pi}<0$), so there is no change of sign, is in red. The region where condition (b) is not satisfied ($Bias_{\pi}^{LR}<GTE_{\pi}$ or $Bias_{\pi}^{CR}<GTE_{\pi}$, respectively), so there is no change of sign, is in blue.
@@ 7.5.2 Gap from Optimality
One important implication of Figure 3 is the impact that the change of sign region has on the net profits of firms who use canonical price experimentation methods. Consider the practical importance of the change of sign region’s borders. The right border of the change of sign region (where purple meets red), represents the optimal price for that particular market balance parameter $\lambda$. This is the point where $GTE_{\pi}=0$. The left border of the change of sign region (where blue meets purple) represents the final price a firm would reach, were it to repeatedly update it’s price in response to canonical A/B experimentation. This is the point where $\widehat{GTE}_{\pi}=0$, or equivalently, where $Bias_{\pi}=GTE_{\pi}$. Note that this price corresponds to the optimum suggested by repeated canonical A/B experiments. This is because, to the right of that price point, the experiment suggests decreasing the price, and to the left, increasing it.
The difference between these two prices (the true optimal price and the experimentally optimized price) can correspond to significant losses in overall profit. In Figure 4, we see the percent of true optimal profit that a firm will achieve if it uses canonical LR (left) or CR (right) experimentation, for a range of market balance scenarios.
Figure 4: Illustration of the gap from optimal profits that canonical LR, and CR A/B experimentation yields, with $v(p)=e^{V-p}$, $V=5$, $c=1$, $\rho=1$, and $\epsilon=1$.
What we find is that the profit achieved from optimization via experimentation can be far below the optimal profit, with the impact worsening in highly demand-constrained markets (for LR experiments) or demand-constrained markets (for CR experiments). Even when markets are reasonably balanced (i.e., $\lambda\approx 1$), both LR and CR experimentation will lead to a profit that is more than 4% lower than optimal. Overall, these findings suggest that firms who use canonical A/B methods for price experimentation might be operating well below their potentially optimal profits, even if they are in well balanced markets. Furthermore, these losses can increase significantly as markets become more imbalanced, getting to $20-40\%$ of optimal.
@@ 8 An Empirical Case Study
In this section, we investigate the feasibility of changes of sign in price experments in a calibrated, real-world setting. In particular, we consider the lodging marketplace Airbnb, using market conditions for a number of cities and market segments estimated by Farronato and Fradkin, (2022). We leverage their estimates to calibrate our change of sign conditions (13) and (14). As we show, the estimates of Farronato and Fradkin, (2022) allow us to obtain interpretable conditions under which a change of sign might occur in a price experiment in each city and market segment. Notably, we demonstrate that for a reasonable range of market conditions, changes of sign would be expected in practice.
We begin by rewriting our change of sign conditions to be more suited to the Farronato and Fradkin, (2022) data that estimates own price elasticities for (city, market) segments. Recall the following two necessary and sufficient conditions for a change of sign to occur for the canonical price experiment we study, cf. (13)-(14):
Condition (a): $GTE_{\pi}(p)\geq 0;$
Condition (b): $Bias_{\pi}(p,q)\geq GTE(p).$
Note that $e_{p}(p)$ is the price elasticity of demand and is defined as $e_{p}(p):=pD^{\prime}(p)/D(p)$. The first of these conditions, Condition (a), ensures that the true global treatment effect of a sufficiently small price increase is positive. In other words, this condition is only satisfied when the firm is currently (locally) “underpricing” relative to the optimal price . We can rewrite this condition using the monopoly pricing optimality condition as follows:
$GTE_{\pi}(p)\geq 0\iff$ $\frac{p}{p-c}+e_{p}(p)\geq 0.$ (28)
The second of these conditions, Condition (b), is satisfied when the interference bias is at least as large as the GTE itself. For the purposes of this exercise, we suppose that $q=1/2$, i.e., that half of all units are treated, and half are left in control. As discussed in Section 6, under this equal allocation experimental design, there holds $D_{1x}(p,p,q)=D_{0y}(p,p,q)$.
In order to rewrite Condition (b), we take advantage of (11):
$Bias_{\pi}(p,q)=(p-c)\left(D_{0y}(p,p,q)+D_{1x}(p,p,q)\right).$
Combining this with the assumption that $q=1/2$, we can rewrite bias as:
$Bias_{\pi}(p,1/2)=2(p-c)D_{0y}(p,p,1/2).$
To proceed, define $e_{0y}(p,p,q)$ as the cross-price elasticity of control demand, i.e., the percentage change in control demand for a one percent increase in the treatment price. Recalling that $D_{0}(p,p,q)=Q_{0}(p,p,q)/(1-q)$, we define $e_{0y}(p,p,q)$ as follows:
$e_{0y}(p,p,q):=\frac{Q_{0y}(p,p,q)p}{Q_{0}(p,p,q)}=\frac{D_{0y}(p,p,q)p}{D_{0}(p,p,q)}.$
We can then write Condition (b) as:
$Bias_{\pi}(p,1/2)\geq GTE_{\pi}(p)\iff$ $2(p-c)D_{0y}(p,p,1/2)\geq D(p)+(p-c)D^{\prime}(p)$
$\iff$ $\frac{2(p-c)}{p}\frac{pD_{0y}(p,p,1/2)}{D(p)}\geq 1+\frac{(p-c)}{p}\frac{pD^{\prime}(p)}{D(p)}$
$\iff$ $\frac{2(p-c)}{p}e_{0y}(p,p,1/2)\geq 1+\frac{(p-c)}{p}e_{p}(p)$
$\iff$ $2e_{0y}(p,p,1/2)\geq\frac{p}{p-c}+e_{p}(p)$ (29)
Note that $D_{0}(p,p,q)=D(p)$ by Assumption 4.2. Now, combining (28)-(29) allows us to write the change of sign conditions as the following chain of inequalities:
$2e_{0y}(p,p,1/2)\geq\frac{p}{p-c}+e_{p}(p)\geq 0$ (30)
Next, we investigate the feasibility of this chain of inequalities holding in real world markets. Observe that in (30) there are four unknown variables: the price $p$, the cost $c$, the own price elasticity of control demand $e_{p}(p)$, and the cross-price elasticity of demand $e_{0y}(p,p,1/2)$. Farronato and Fradkin, (2022) reports estimates of the costs and own price elasticities of demand for Airbnb market segments (economy, midscale, upscale, luxury) in 10 cities around the United States. This leaves us with two missing variables: the price and the cross-price elasticity.
It is important to note that the cross-price elasticity, $e_{0y}(p,p,1/2)$, has a different interpretations in a two-sided market like Airbnb when running LR and CR experiments. In an LR experiment, $e_{0y}(p,p,1/2)$ quantifies the change in demand for control listings when the price of treated listings is raised. The experiment effectively creates two categories of products with higher (treatment) and lower (control) prices. If treatment raises the price, then control demand rises because price sensitive customers move away from the higher priced treatment listings, towards the control listings. The cross-price elasticity captures this effect, exactly as usual in demand modeling.
In a CR experiment, $e_{0y}(p,p,1/2)$ has a different interpretation. Now, $e_{0y}(p,p,1/2)$ quantifies the degree to which demand from control customers changes when the price is raised for treatment customers. The mechanism by which this occurs depends on an availability effect: if the treatment raises the price, then as demand from treatment customers falls, product availability increases, which in turn positively impacts control demand. To demonstrate this more formally, with a slight abuse of notation, let $D(p,s)$ represent demand when price is $p$, and available inventory is $s$ (out of total inventory $\rho$). If we run a CR experiment on a fraction $q$ of customers, the following balance equation must hold:
$(1-q)D(p_{0},s)+qD(p_{1},s)=\rho-s$
This equation implicitly defines $s$ in terms of $p_{0},p_{1}$, and $q$. Indeed, observe that this a more general version of the CR steady-state condition given for the logit demand model in (25). If we now differentiate with respect to $p_{1}$, we obtain:
$(1-q)D_{s}(p_{0},s)\frac{\partial s}{\partial p_{1}}+q\left[D_{p}(p_{1},s)+D_{s}(p_{1},s)\frac{\partial s}{\partial p_{1}}\right]=-\frac{\partial s}{\partial p_{1}}$
Setting $p_{1}=p_{0}$ gives:
$D_{s}(p_{0},s)\frac{\partial s}{\partial p_{1}}=-\frac{\partial s}{\partial p_{1}}-qD_{p}(p_{0},s)$ (31)
In (31), the left hand side represents the change in rescaled control demand as the treatment price is raised; this is $D_{0y}(p,p,q)$ for CR experiments. The right hand side shows that this quantity depends on two effects: a direct price effect on demand, $D_{p}(p_{0},s)$; but also the impact of the price change on equilibrium availability, $\partial s/\partial p_{1}$. Measuring this quantity requires empirical calibration of the change in supply rationing when prices are changed for a subset of customers. With published estimates such as those in Farronato and Fradkin, (2022), it is difficult to estimate any reasonable range for this availability effect.
We therefore focus on LR experimentation in our presentation, rather than CR experimentation. Although $e_{0y}(p,p,1/2)$ is not observed, it is possible to consider a reasonable range for this cross elasticity, by comparing it to the underlying own price elasticity of demand in each market segment as reported in Farronato and Fradkin, (2022). As we have discussed above, LR price experimentation is the industry standard. Therefore this approach also captures the regime of most interest to practitioners.
Returning now to the chain of inequalities which produce a sign flip (30), we can use the data reported in Farronato and Fradkin, (2022) to plot ranges of prices and cross-price elasticities for which we would expect to see changes of sign in canonical LR price experiments. For price, we consider a range beginning at cost, and ending slightly above the optimal monopoly price implied by $\hat{e}_{p}$ and $\hat{c}$. For the cross-price elasticity we plot beginning at 0, and end at $2\times-\hat{e}_{p}$. We choose this range of cross-price elasticities because in general we expect the cross-price elasticity to be commensurate with the magnitude of the own price elasticity in LR experiments. Since the true cross-price elasticity is unknowable given our data, we select this range to be conservative. We should note that while it is possible to have a cross-price elasticity that exceeds $2\times-\hat{e}_{p}$, in these settings the interference bias will be so large that a sign flip is even more likely if the firm is underpricing.
Figure 5: Change of sign region for price and cross-price elasticity in the Airbnb Boston Economy Segment. Farronato and Fradkin, (2022) estimates $\hat{c}=\mathdollar 81.4$, and $\hat{e}_{p}(p):-2.33$.
Figure 5 shows the change of sign region plot for the economy segment of Airbnb listings in Boston. (See Appendix 11 for similar plots for the other market segments, as well as the other cities we consider; the results are similar to Figure 5.) In the blue region the bias is too low to cause a change of sign (Condition (b) is violated), while in the red region the current price is already too high (Condition (a) is violated). For parameters in the purple region, our analysis suggests that changes of sign will occur. It is important to note that since we do not observe the market price or the cross-price elasticity of demand, the estimates in Farronato and Fradkin, (2022) are insufficient to identify the exact state of the world on this plot.
Nevertheless, one can reasonably conclude from Figure 5 that changes of sign are plausible. In particular, if listings are even mildly underpriced, and the cross-price elasticity of control demand is comparable to the magnitude of the own-price elasticity of demand (i.e., approximately $2$), then our results show that changes of sign would arise: the experiment will incorrectly suggest that the current price is too high, and the experimenter will be encouraged to move the price in the wrong direction. We view these results as compelling evidence that changes of sign are a phenomenon that can have significant consequence for practitioners.